OpenStudy (anonymous):

n/12= 300 n = ? (n/12 is a fraction)

sam (.sam.):

multiply 300 by 12

OpenStudy (anonymous):

is that is?

sam (.sam.):

3600

Directrix (directrix):

n/12= 300 n/12 = 300/1 Cross multiply n(1) = 12(300) n = 3600